# Interaction-induced transition in quantum many-body detection probability

### 06 Mar 2024 v2

@@ Abstract
.1 Necessary and sufficient condition for unit spectral radius of $\hat{M}_{Q}(\tau)$ at any finite value of $\tau$.
@@
$\hat{M}_{Q}(\tau)=\hat{Q}\Big{[}\hat{U}_{Q}(\tau)+\sum_{p=2}^{\infty}\frac{(-i)^{p}\tau^{p}}{p!}\left\{\hat{H}^{p}-\left(\hat{Q}\hat{H}\hat{Q}\right)^{p}\right\}\Big{]}\hat{Q}$ (6)
@@
$\hat{M}_{Q}(\tau)\simeq\hat{Q}\hat{U}_{Q}(\tau)\hat{Q}-\frac{\tau^{2}}{2}\left(\hat{Q}\hat{H}^{2}\hat{Q}-\hat{Q}\hat{H}\hat{Q}\hat{H}\hat{Q}\right).$ (7)
@@
$e^{-\lambda_{m}(\tau)+i\theta_{m}(\tau)}\simeq e^{-i\tau E^{Q}_{m}}-\frac{\tau^{2}}{2}\left[\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}\right].$ (8)
@@
$\lambda_{m}(\tau)=\tau^{2}\left[\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}\right]+O(\tau^{3}).$ (9)
@@
Next, since $\lambda_{m}(\tau)\geq 0$, we have $\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}\geq 0$. Consequently, for finite but small $\tau$, the signal is always eventually detected, unless, for some choice of $m$, we have $\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}=0$. The later condition can only happen if $|E^{Q}_{m}\rangle$ is an eigenvector of $\hat{H}$ with eigenvalue $E^{Q}_{m}$. Therefore, this eigenvector of $\hat{H}$ must belong completely to the $\hat{Q}$ subspace. However, if this is the case, by multiplying Eq.(6) with $|E^{Q}_{m}\rangle$ from right, we directly see that it is an eigenvector of $\hat{M}_{Q}$, the corresponding eigenvalue having $\lambda_{m}(\tau)=0$ irrespective of the value of $\tau$. Thus, some eigenvectors of $\hat{H}$ belonging completely to the $\hat{Q}$ subspace is a necessary condition for unit spectral radius of $\hat{M}_{Q}(\tau)$ for any finite value of $\tau$.
.2 System and initial state
@@
$\hat{H}=-\sum_{\ell=-N/2+1}^{N/2-1}\left[\frac{J}{2}\left(\hat{c}_{\ell}^{\dagger}\hat{c}_{\ell+1}+\hat{c}_{\ell+1}^{\dagger}\hat{c}_{\ell}\right)+\Delta\hat{n}_{\ell}\hat{n}_{\ell+1}\right]+\varepsilon_{0}\hat{n}_{0},$ (10)
@@
$\hat{H}_{1}=-\frac{J}{2}\left(\hat{c}_{1}^{\dagger}\hat{c}_{0}+\hat{c}_{0}^{\dagger}\hat{c}_{1}\right),~{}~{}\hat{H}_{0}=\hat{H}-\hat{H}_{1}.$ (11)
@@
For single-trajectory simulation— For this purpose, we directly simulate the action of the two detectors on the dynamics as follows. The detectors makes stroboscopic measurements in steps of time $\tau$. Just before the $n$th stroboscopic measurement, let the state of the system be $|\psi_{b}(n\tau)\rangle$. For detectors placed at sites $p$ and $q$, we calculate the expectation values
$\langle\hat{n}_{p}(n\tau)\rangle=\langle\psi_{b}(n\tau)|\hat{n}_{p}|\psi_{b}(n\tau)\rangle$
$\langle\hat{n}_{q}(n\tau)\rangle=\langle\psi_{b}(n\tau)|\hat{n}_{q}|\psi_{b}(n\tau)\rangle.$ (13)
These give the probabilities of detecting a particle at $p$ and $q$ sites independently. To simulate the action of the detectors, we choose two independent random numbers, $r_{p}$, $r_{q}$, from a uniform distribution between 00 and $1$, and obtain $|\psi_{f}(n\tau)\rangle$ as follows,
${\rm if}~{}~{}r_{p}\leq\langle\hat{n}_{p}(n\tau)\rangle~{}~{}{and}~{}~{}r_{q}\leq\langle\hat{n}_{q}(n\tau)\rangle:$
@@
${\rm if}~{}~{}r_{p}\leq\langle\hat{n}_{p}(n\tau)\rangle~{}~{}{and}~{}~{}r_{q}>\langle\hat{n}_{q}(n\tau)\rangle:$
$\qquad|\psi_{f}(n\tau)\rangle=\hat{n}_{p}(1-\hat{n}_{q})|\psi_{b}(n\tau)\rangle,$
${\rm if}~{}~{}r_{p}>\langle\hat{n}_{p}(n\tau)\rangle~{}~{}{and}~{}~{}r_{q}\leq\langle\hat{n}_{q}(n\tau)\rangle:$
$\qquad|\psi_{f}(n\tau)\rangle=(1-\hat{n}_{p})\hat{n}_{q}|\psi_{b}(n\tau)\rangle,$
${\rm if}~{}~{}r_{p}>\langle\hat{n}_{p}(n\tau)\rangle~{}~{}{and}~{}~{}r_{q}>\langle\hat{n}_{q}(n\tau)\rangle:$
$\qquad|\psi_{f}(n\tau)\rangle=(1-\hat{n}_{p})(1-\hat{n}_{q})|\psi_{b}(n\tau)\rangle.$ (14)
@@
Here we show the plots of dynamics of the system, in absence of any measurements, starting from the two different initial states corresponding to $E=E^{Q_{0}}_{1}$, $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. The plots are shown in Fig. 3.
@@
In Fig. 3(a), when $E=E^{Q_{0}}_{1}$, we see that for $\Delta<J$, $N_{R}(t)$ (where we now explicitly write the time argument) increases with $t$ towards $N_{R}(t)\sim N/2$. For $\Delta>J$, $N_{R}(t)$ remains close to $N_{R}(t)\sim 1$, over the entire time range considered. When $E=E^{Q_{0}}_{\alpha_{\rm mid}}$, $N_{R}(t)$ increases towards $N/2$ for all values of $\Delta$, an shown in Fig. 3(b). So, in the domain wall picture, we clearly see that the domain wall does not melt for $\Delta>J$ and $E=E^{Q_{0}}_{1}$, while in other cases, it melts, as expected in non-integrable systems.
@@
Figure 3: (a) Dynamics of number of particles in right half, $N_{R}(t)$, for various values of $\Delta$, when starting from an initial state of the form Eq.(12) with $E=E_{1}^{Q_{0}}$. (b) The same with $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. (c) Dynamics of occupation at site $p=3$, for various values of $\Delta$, when when initial state is chosen with $E=E^{Q_{0}}_{1}$ in Eq.(4) of main text. (d) The same with $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. Other parameters, $N=22$, $\epsilon_{0}=0.5J$, $\sigma=0.1J$.
@@
In Fig. 3(c), we plot the dynamics of occupation at site $p$, $\langle\hat{n}_{p}(t)\rangle$, with $p=3$, for various values of $\Delta$, starting with the initial state corresponding to $E=E^{Q_{0}}_{1}$. The occupation approaches $\langle\hat{n}_{p}(t)\rangle\sim 0.5$ for $\Delta<J$, while it shows oscillations about a much smaller value for $\Delta>J$. The important point to note here is that even for $\Delta>J$, where the domain wall does not melt, the dynamics is not completely frozen. Rather, it is restricted to a small subspace. In Fig. 3(d), we show similar plots starting with the initial state corresponding to $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. In this case, the occupation approaches $\langle\hat{n}_{p}(t)\rangle\sim 0.5$ for all values of $\Delta$ considered here.
.5 Single detection after time $n\tau$
@@
Figure 4: Panel (a) shows plot of the expectation value $\langle\hat{n}_{p}\hat{n}_{q}\rangle$ at time $2000J^{-1}$ versus $\Delta/J$ for various system sizes, when starting from an initial state of the form Eq.(12) with $E=E_{1}^{Q_{0}}$. The black dots shows the same for $N=26$ when starting from an initial state of the form Eq.(12) with $E=E_{\alpha_{\rm mid}}^{Q_{0}}$. Panel (b) shows the same plot with y-axis in log scale. This expectation value gives the probability of getting simultaneous clicks on two detectors at sites $p$ and $q$ in a single measurement done at time $n\tau$. Parameters: $p=3$, $q=5$, $\varepsilon_{0}=0.5J$, $\sigma=0.1J$.
@@
Instead of stroboscopic measurements at steps of time $\tau$, we can also ask what is the probability of simultaneous detection at sites $p$ and $q$ in a single measurement after a given time. This is given by the expectation value $\langle\hat{n}_{p}\hat{n}_{q}\rangle$ calculated at that time. In Fig. 4, we show plots of this expectation value at time $2000J^{-1}$ as a function of $\Delta/J$. When starting with an initial state of the form Eq.(12) with $E=E_{1}^{Q_{0}}$, we see that $\langle\hat{n}_{p}\hat{n}_{q}\rangle$ shows a smooth crossover from a finite value to a small value on going across $\Delta/J=1$. Clearly, the probability of simultaneous detection on both sites, i.e, of detecting the ‘signal’, is not close to $1$ even when $\Delta/J<1$, within the chosen parameter regime. There is still considerable probability of not detecting the signal ($>70\%$ in the chosen parameter regime). This is very different from stroboscopic measurements where the no-detection probability was $<10^{-12}$ for $\Delta/J<0.9$, and sharply rises to $\approx 1$ for $\Delta/J=1.1$. Thus, it is clear that backaction of stroboscopic measurements makes the transition much sharper.
@@
Figure 5: The figure shows that value of $\Delta_{*}/J$ as a function of system size $N$. Parameters, $E=E^{Q_{0}}_{1}$, $\epsilon=10^{-5}$, $n=1000$, $\varepsilon_{0}=0.5J$, $\sigma=0.1J$.
.6 Behavior versus $N$
@@
We now show the effect of changing system sizes on the sharp transition in $R_{n}$. For the initial state of the form Eq.(12) with $E=E_{1}^{Q_{0}}$, we define $\Delta_{*}/J$ has the smallest value of $\Delta/J$ for which $R_{n}>\epsilon$, where $\epsilon$ is some chosen tolerance. In Fig. 5, we show the plot for $n=1000$ and $\epsilon=10^{-5}$. We see that the value of $\Delta_{*}/J$ changes only a little on nearly doubling the system size, and seems to saturate for larger system size. This gives further numerical evidence for the sharpness of the transition in $R_{n}$.
@@
Figure 6: a) Plot of $\lambda_{1}(\tau)$ with $\Delta$, for various system sizes, and for $\tau=4J^{-1}$. (b) Plot of no-detection probability, $R_{n}$, with $\Delta$ after $n=1000$ steps, for the same system sizes, and for $\tau=4J^{-1}$. Other parameters: $E=E^{Q_{0}}_{1}$, $\varepsilon_{0}=0.5J$, $\sigma=0.1J$.
.7 Different choice of $\tau$
@@
Our analytical understanding, as well as our numerical results on domain-wall melting, shows that the sharp transition in QMBDP is independent of the choice of $\tau$. In Fig. 6 we show plots of $\lambda_{1}(\tau)$ versus $\Delta$, and no-detection probability $R_{n}$ for $n=1000$, and $\tau=4J^{-1}$. This value of $\tau$ is different from that of the plots in the main text, where $\tau=2J^{-1}$. We still see the same behavior, no-detection probability changes from $R_{n}\sim 10^{-12}$ to $R_{n}\sim 1$ on changing $\Delta/J$ from $0.9$ to $1.1$.
@@
Figure 7: Comparison of single trajectory plots showing the transition in cases where the two detectors are placed at (a) $p=3$, $q=6$, (b) $p=3$, $q=5$. Here $C$ is the number of simultaneous clicks in the two detectors after $n=1024$ single-shot stroboscopic measurements. Three trajectories, i.e, simulation of three runs of the experiment, are shown for each case. Other parameters, $N=20$, $\epsilon_{0}=0.5J$, $\tau=2J^{-1}$.
.8 Different position of detectors
@@
All numerical results for the transition in QMBDP presented till now are for the case where the detectors are placed at sites $p=3$ and $q=5$. Our analytical understanding, as well as our numerical results on domain-wall melting show that they remain unchanged for any positions of the detectors, at a finite distance from the middle. To explicitly show this, in Fig. 7, we plot the number of simultaneous clicks in single trajectory simulations of up to $n=1024$ stroboscopic measurements for the cases with detectors at $p=3,q=6$, and $p=3,q=5$, as function of $\Delta/J$. We clearly see identical behavior, with transition at $\Delta/J=1$.

### 20 Mar 2024 v3

@@ Abstract
S1 Necessary and sufficient condition for unit spectral radius of $\hat{M}_{Q}(\tau)$ at any finite value of $\tau$.
@@
$\hat{M}_{Q}(\tau)=\hat{Q}\Big{[}\hat{U}_{Q}(\tau)+\sum_{p=2}^{\infty}\frac{(-i)^{p}\tau^{p}}{p!}\left\{\hat{H}^{p}-\left(\hat{Q}\hat{H}\hat{Q}\right)^{p}\right\}\Big{]}\hat{Q}$ (S2)
@@
$\hat{M}_{Q}(\tau)\simeq\hat{Q}\hat{U}_{Q}(\tau)\hat{Q}-\frac{\tau^{2}}{2}\left(\hat{Q}\hat{H}^{2}\hat{Q}-\hat{Q}\hat{H}\hat{Q}\hat{H}\hat{Q}\right).$ (S3)
@@
$e^{-\lambda_{m}(\tau)+i\theta_{m}(\tau)}\simeq e^{-i\tau E^{Q}_{m}}-\frac{\tau^{2}}{2}\left[\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}\right].$ (S4)
@@
$\lambda_{m}(\tau)=\tau^{2}\left[\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}\right]+O(\tau^{3}).$ (S5)
@@
Next, since $\lambda_{m}(\tau)\geq 0$, we have $\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}\geq 0$. Consequently, for finite but small $\tau$, the signal is always eventually detected, unless, for some choice of $m$, we have $\langle E^{Q}_{m}|\hat{H}^{2}|E^{Q}_{m}\rangle-\left(E^{Q}_{m}\right)^{2}=0$. The later condition can only happen if $|E^{Q}_{m}\rangle$ is an eigenvector of $\hat{H}$ with eigenvalue $E^{Q}_{m}$. Therefore, this eigenvector of $\hat{H}$ must belong completely to the $\hat{Q}$ subspace. However, if this is the case, by multiplying Eq.(S2) with $|E^{Q}_{m}\rangle$ from right, we directly see that it is an eigenvector of $\hat{M}_{Q}$, the corresponding eigenvalue having $\lambda_{m}(\tau)=0$ irrespective of the value of $\tau$. Thus, some eigenvectors of $\hat{H}$ belonging completely to the $\hat{Q}$ subspace is a necessary condition for unit spectral radius of $\hat{M}_{Q}(\tau)$ for any finite value of $\tau$.
S2 System and initial state
@@
$\hat{H}=-\sum_{\ell=-N/2+1}^{N/2-1}\left[\frac{J}{2}\left(\hat{c}_{\ell}^{\dagger}\hat{c}_{\ell+1}+\hat{c}_{\ell+1}^{\dagger}\hat{c}_{\ell}\right)+\Delta\hat{n}_{\ell}\hat{n}_{\ell+1}\right]+\varepsilon_{0}\hat{n}_{0},$ (S6)
@@
$\hat{H}_{1}=-\frac{J}{2}\left(\hat{c}_{1}^{\dagger}\hat{c}_{0}+\hat{c}_{0}^{\dagger}\hat{c}_{1}\right),~{}~{}\hat{H}_{0}=\hat{H}-\hat{H}_{1}.$ (S7)
@@
$\left|\psi(0)\right\rangle\propto{\rm exp}\left[-\left(\frac{\hat{H}_{0}-E}{\sigma}\right)^{2}\right]\left|\psi_{\rm rand}\right\rangle_{N_{R}=1},$ (S8)
@@
For single-trajectory simulation— For this purpose, we directly simulate the action of the two detectors on the dynamics as follows. The detectors makes stroboscopic measurements in steps of time $\tau$. Just before the $n$th stroboscopic measurement, let the state of the system be $|\psi_{b}(n\tau)\rangle$. For detectors placed at sites $p$ and $q$, we calculate the expectation value
$\langle\hat{n}_{p}(n\tau)\hat{n}_{q}(n\tau)\rangle=\langle\psi_{b}(n\tau)|\hat{n}_{p}\hat{n}_{q}|\psi_{b}(n\tau)\rangle$ (S9)
This gives the probability of simultaneous detection at $p$ and $q$ sites. To simulate the action of the detectors, we choose an independent random numbers, $r$ from a uniform distribution between 00 and $1$, and obtain $|\psi_{f}(n\tau)\rangle$ as follows,
${\rm if}~{}~{}r\leq\langle\hat{n}_{p}(n\tau)\hat{n}_{q}(n\tau)\rangle:$
@@
${\rm else}:$
$\qquad|\psi_{f}(n\tau)\rangle=|\psi_{b}(n\tau)\rangle-\hat{n}_{p}\hat{n}_{q}|\psi_{b}(n\tau)\rangle$ (S10)
@@
Here we show the plots of dynamics of the system, in absence of any measurements, starting from the two different initial states corresponding to $E=E^{Q_{0}}_{1}$, $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. The plots are shown in Fig. S1.
@@
In Fig. S1(a), when $E=E^{Q_{0}}_{1}$, we see that for $\Delta<J$, $N_{R}(t)$ (where we now explicitly write the time argument) increases with $t$ towards $N_{R}(t)\sim N/2$. For $\Delta>J$, $N_{R}(t)$ remains close to $N_{R}(t)\sim 1$, over the entire time range considered. When $E=E^{Q_{0}}_{\alpha_{\rm mid}}$, $N_{R}(t)$ increases towards $N/2$ for all values of $\Delta$, an shown in Fig. S1(b). So, in the domain wall picture, we clearly see that the domain wall does not melt for $\Delta>J$ and $E=E^{Q_{0}}_{1}$, while in other cases, it melts, as expected in non-integrable systems.
@@
Figure S1: (a) Dynamics of number of particles in right half, $N_{R}(t)$, for various values of $\Delta$, when starting from an initial state of the form Eq.(S8) with $E=E_{1}^{Q_{0}}$. (b) The same with $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. (c) Dynamics of occupation at site $p=3$, for various values of $\Delta$, when when initial state is chosen with $E=E^{Q_{0}}_{1}$ in Eq.(4) of main text. (d) The same with $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. Other parameters, $N=22$, $\epsilon_{0}=0.5J$, $\sigma=0.1J$.
@@
In Fig. S1(c), we plot the dynamics of occupation at site $p$, $\langle\hat{n}_{p}(t)\rangle$, with $p=3$, for various values of $\Delta$, starting with the initial state corresponding to $E=E^{Q_{0}}_{1}$. The occupation approaches $\langle\hat{n}_{p}(t)\rangle\sim 0.5$ for $\Delta<J$, while it shows oscillations about a much smaller value for $\Delta>J$. The important point to note here is that even for $\Delta>J$, where the domain wall does not melt, the dynamics is not completely frozen. Rather, it is restricted to a small subspace. In Fig. S1(d), we show similar plots starting with the initial state corresponding to $E=E^{Q_{0}}_{\alpha_{\rm mid}}$. In this case, the occupation approaches $\langle\hat{n}_{p}(t)\rangle\sim 0.5$ for all values of $\Delta$ considered here.
S5 Single detection after time $n\tau$
@@
Figure S2: Panel (a) shows plot of the expectation value $\langle\hat{n}_{p}\hat{n}_{q}\rangle$ at time $2000J^{-1}$ versus $\Delta/J$ for various system sizes, when starting from an initial state of the form Eq.(S8) with $E=E_{1}^{Q_{0}}$. The black dots shows the same for $N=26$ when starting from an initial state of the form Eq.(S8) with $E=E_{\alpha_{\rm mid}}^{Q_{0}}$. Panel (b) shows the same plot with y-axis in log scale. This expectation value gives the probability of getting simultaneous clicks on two detectors at sites $p$ and $q$ in a single measurement done at time $n\tau$. Parameters: $p=3$, $q=5$, $\varepsilon_{0}=0.5J$, $\sigma=0.1J$.
@@
Instead of stroboscopic measurements at steps of time $\tau$, we can also ask what is the probability of simultaneous detection at sites $p$ and $q$ in a single measurement after a given time. This is given by the expectation value $\langle\hat{n}_{p}\hat{n}_{q}\rangle$ calculated at that time. In Fig. S2, we show plots of this expectation value at time $2000J^{-1}$ as a function of $\Delta/J$. When starting with an initial state of the form Eq.(S8) with $E=E_{1}^{Q_{0}}$, we see that $\langle\hat{n}_{p}\hat{n}_{q}\rangle$ shows a smooth crossover from a finite value to a small value on going across $\Delta/J=1$. Clearly, the probability of simultaneous detection on both sites, i.e, of detecting the ‘signal’, is not close to $1$ even when $\Delta/J<1$, within the chosen parameter regime. There is still considerable probability of not detecting the signal ($>70\%$ in the chosen parameter regime). This is very different from stroboscopic measurements where the no-detection probability was $<10^{-12}$ for $\Delta/J<0.9$, and sharply rises to $\approx 1$ for $\Delta/J=1.1$. Thus, it is clear that backaction of stroboscopic measurements makes the transition much sharper.
@@
Figure S3: The figure shows that value of $\Delta_{*}/J$ as a function of system size $N$. Parameters, $E=E^{Q_{0}}_{1}$, $\epsilon=10^{-5}$, $n=1000$, $\varepsilon_{0}=0.5J$, $\sigma=0.1J$.
S6 Behavior versus $N$
@@
We now show the effect of changing system sizes on the sharp transition in $R_{n}$. For the initial state of the form Eq.(S8) with $E=E_{1}^{Q_{0}}$, we define $\Delta_{*}/J$ has the smallest value of $\Delta/J$ for which $R_{n}>\epsilon$, where $\epsilon$ is some chosen tolerance. In Fig. S3, we show the plot for $n=1000$ and $\epsilon=10^{-5}$. We see that the value of $\Delta_{*}/J$ changes only a little on nearly doubling the system size, and seems to saturate for larger system size. This gives further numerical evidence for the sharpness of the transition in $R_{n}$.
@@
Figure S4: a) Plot of $\lambda_{1}(\tau)$ with $\Delta$, for various system sizes, and for $\tau=4J^{-1}$. (b) Plot of no-detection probability, $R_{n}$, with $\Delta$ after $n=1000$ steps, for the same system sizes, and for $\tau=4J^{-1}$. Other parameters: $E=E^{Q_{0}}_{1}$, $\varepsilon_{0}=0.5J$, $\sigma=0.1J$.
S7 Different choice of $\tau$
@@
Our analytical understanding, as well as our numerical results on domain-wall melting, shows that the sharp transition in QMBDP is independent of the choice of $\tau$. In Fig. S4 we show plots of $\lambda_{1}(\tau)$ versus $\Delta$, and no-detection probability $R_{n}$ for $n=1000$, and $\tau=4J^{-1}$. This value of $\tau$ is different from that of the plots in the main text, where $\tau=2J^{-1}$. We still see the same behavior, no-detection probability changes from $R_{n}\sim 10^{-12}$ to $R_{n}\sim 1$ on changing $\Delta/J$ from $0.9$ to $1.1$.
@@
Figure S5: Comparison of single trajectory plots showing the transition in cases where the two detectors are placed at (a) $p=3$, $q=6$, (b) $p=3$, $q=5$. Here $C$ is the number of simultaneous clicks in the two detectors after $n=1024$ single-shot stroboscopic measurements. Three trajectories, i.e, simulation of three runs of the experiment, are shown for each case. Other parameters, $N=20$, $\epsilon_{0}=0.5J$, $\tau=2J^{-1}$.
S8 Different position of detectors
@@
All numerical results for the transition in QMBDP presented till now are for the case where the detectors are placed at sites $p=3$ and $q=5$. Our analytical understanding, as well as our numerical results on domain-wall melting show that they remain unchanged for any positions of the detectors, at a finite distance from the middle. To explicitly show this, in Fig. S5, we plot the number of simultaneous clicks in single trajectory simulations of up to $n=1024$ stroboscopic measurements for the cases with detectors at $p=3,q=6$, and $p=3,q=5$, as function of $\Delta/J$. We clearly see identical behavior, with transition at $\Delta/J=1$.